The rate constant for a reaction at 40 degree C is exactly 3 times that at 20 degree C. Calculate the activation energy for the reaction. R=8.314J/Kmol

K=Ae^(-Ea/RT)
Answer #1

Ams initial temperature (T) = 20°c = 20+ 273 = 293 K final temperature (72) = 4oc 40 + 273 = 313 K Rate Constant Ratio A = 3